# Driven-dissipative turbulence in exciton-polariton quantum fluids

### v2

@@ IV Conclusion
In this work, we studied the 2D quantum turbulence with a focus on the inclusion of important specialties inherent to quantum fluids of exciton-polaritons in semiconductor microcavities. The first exciton-polariton feature complication the turbulence studies in a non-trivial implementation of the spoon-stirring vortex injection for turbulence development. However, we demonstrate robustly that one can effectively generate turbulence by direct imprinting the wave function with homogeneous magnitude and phase formed of tiles with randomly oriented plane waves. In experimental implementation of this scheme, one can use SLM with quasi-resonant laser excitation. The size of the tiles should be comparable to the desired distance between the layers. The created pattern automatically creates phase singularities, which then evolve into quantized vortices. After a certain free evolution period, the resulting statistics of the vortex and incompressible kinetic energy spectra are very similar to those of the conventional spoon-stirring technique. For clustered vortex only, we see the -5/3 exponent in the IKE spectrum that is likely related to Kolmogorov law. The slightly higher clusterization for Tile-8 scheme is visible in the correlation function for the signs of nearest neighbors. The randomness of the imprinted wave vector directions in Tile-8 provides the inherent stochasticity of quantum turbulence and is useful for averaging both for numerical simulations and for experimental realizations. At the same time, the imprinting scheme can allow one to collect the signal in repetitive experiments. Therefore, we conclude that the Tile-8 imprinting scheme is a valid alternative to conventional spoon stirring to study quantum turbulence in quantum fluids. Excitation with pixel random phase is also capable to generate vortices, however, efficiency of generation is significantly lower in this case. Moreover, creation of turbulent-like flows and clusters is also less effective for pixel schemes, which means that the tile schemes posses important features not occurring in more random approach.
@@
Secondly, we show that quantum turbulence can be robust with respect to polariton decay. The statistical quantities like the total number of vortices, the fraction of single and clustered vortices, fraction of vortices belonging to the dipoles evolve smoothly with depletion of the polariton density. The spectral characteristics of the quantum fluid also evolve smoothly. For sufficiently slow decay rate (high lifetimes), IKE spectra do not depend on the decay rate itself, but only on the current density. In general, the fingerprints of turbulence in statistics and IKE spectra remain until approximately one-order decrease of the quantum fluid density. With respect to observation of the evaporative heating effect, one should track the system for at least several nanoseconds for the typical polariton mass and interaction strength, which imposes the respective limitation on the lifetime. Finally, we justify that gain/loss fluctuations do not affect the observation of turbulence under certain conditions on the involved time scales, highlighting the potential utility in quantum turbulence experiments of an experimental scheme of the laser periodically operating above and below the condensation threshold with characteristic period less than polariton lifetime.

### v3

@@ IV Conclusion
In this work, we studied the 2D quantum turbulence with a focus on the inclusion of important specialties inherent to quantum fluids of exciton-polaritons in semiconductor microcavities. The first exciton-polariton feature complication the turbulence studies in a non-trivial implementation of the spoon-stirring vortex injection for turbulence development. However, we demonstrate robustly that one can effectively generate turbulence by direct imprinting the wave function with homogeneous magnitude and phase formed of tiles with randomly oriented plane waves. In experimental implementation of this scheme, one can use SLM with quasi-resonant laser excitation. The size of the tiles should be comparable to the desired distance between the layers. The created pattern automatically creates phase singularities, which then evolve into quantized vortices. After a certain free evolution period in the preparatory phase, the resulting statistics of the vortex and incompressible kinetic energy spectra are very similar to those of the conventional spoon-stirring technique. For clustered vortex only, we see the -5/3 exponent in the IKE spectrum that is related to Kolmogorov scaling law and to the fractal structures of clustered vortices. In addition to clear -5/3 exponent above the forcing scale [40], existence of the associated inverse energy cascade was evidenced by the explicit tracing the time-dynamics of energy distribution [37]. The slightly higher clusterization for Tile-8 scheme is visible in the correlation function for the signs of nearest neighbors. The randomness of the imprinted wave vector directions in Tile-8 provides the inherent stochasticity of quantum turbulence and is useful for averaging both for numerical simulations and for experimental realizations. At the same time, the imprinting scheme can allow one to collect the signal in repetitive experiments. Therefore, we conclude that the Tile-8 imprinting scheme is a valid alternative to conventional spoon stirring to study quantum turbulence in quantum fluids. Excitation with pixel random phase is also capable to generate vortices, however, efficiency of generation is significantly lower in this case. Moreover, creation of turbulent-like flows and clusters is also less effective for pixel schemes, which means that the tile schemes posses important features not occurring in more random approach. It is worth mentioning that, for the condensates in a parabolic traps, the spectral representation of the wave function can be employed for the effective solution of GPE [41]. In that case, it is natural to perform an analysis of the mode occupation number revealing the distribution reminiscent of the wave-turbulent state for the dipolar exciton condensate [42]. However, for strong turbulence with multiple quantized vortices, the chosen numerical scheme is preferable as well as the metrics under analysis: vortex statistics and IKE spectra. In terms of rotational motion scale, the actual turbulent regime lies between that less chaotic wave turbulence-like regime and the regime of Einstein-Bose condensed vortices [7].
@@
Secondly, we show that quantum turbulence can be robust with respect to polariton decay. The statistical quantities like the total number of vortices, the fraction of single and clustered vortices, fraction of vortices belonging to the dipoles evolve smoothly with depletion of the polariton density. The spectral characteristics of the quantum fluid also evolve smoothly. For sufficiently slow decay rate (high lifetimes), IKE spectra do not depend on the decay rate itself, but only on the current density. In general, the fingerprints of turbulence in statistics and IKE spectra remain until approximately one-order decrease of the quantum fluid density. With respect to observation of the evaporative heating effect, one should track the system for at least several nanoseconds for the typical polariton mass and interaction strength, which imposes the respective limitation on the lifetime. In case of enough long-living polaritions comparable to that in Refs. [43, 12] or higher, no need in the additional incoherent drive will be needed, thus excluding the exciton reservoir effects. Finally, we justify that gain/loss fluctuations do not affect the observation of turbulence under certain conditions on the involved time scales, highlighting the potential utility in quantum turbulence experiments of an experimental scheme of the laser periodically operating above and below the condensation threshold with characteristic period less than polariton lifetime.
@@ Acknowledgements
Appendix A NUMERICAL SIMULATIONS
The investigation the efficiency of the wave function imprinting schemes as turbulence excitation techniques in comparison with spoon-stirring procedure is based on numerical simulation of the following driven-dissipative Gross-Pitaevskii equation:
The considered polariton system is characterized as follows: polariton mass (bare electron mass g), interaction constant . The time integration is solved by means of third order Adam-Bashforth method, with time step ps, total duration ns and a square mesh grid of size m. For the calculation of the Laplace operator (kinetic energy term), we used the Fourier transform with massive parallelization of matrix operations provided by GPU in PyTorch.
Within this most general formulation, the dynamical equation above enables to perform analysis for both “standard” stirring techniques (widely used in atomic Bose-Einstein condensates) based for a time-dependent rotating potential and phase-pattern imprinting schemes with stationary potential, . The confining circular hard wall potential of 480 m in diameter, , was normally used. Moreover, all the considered simulations consist of an initial preparatory stage, especially required by the spoon stirring method, and an equally-long analysis stage to investigate quantum turbulence.
The simulation of spoon-stirring was carried out by a 10 m propagating potential with a diameter and rotating at 1.3 m/ps with an orbit radius during the preparatory phase (3 complete spoon rotations). Considering a depth and rotational period ps, the total potential is given by a stirring potential defined as , with Heaviside step function and
The time-modulation amplitude function is considered to smoothly stir on time the condensate during the whole stirring interval, avoiding to generate excessive density noise. For such a case the initial configuration is is an homogeneous state with density . For averaging, the spoon-stirring configurations, we have changed the orbit radius (within 1%).
As for the imprinting methods, two different families of phase pattern were considered: Tile and Pixel Imprinting schemes. In the first case, we divide the wave function into the 8x8 grid of the square tiles of size 64 m and fill each of them with a randomly oriented plane wave with wave length 12 m (Tile-12), 8 m (Tile-8) or 4 m (Tile-4). For the sake of completeness, this imprint method was also studied also 32x32, 16x16 and 4x4 grids with square size respectively equal to 16 m, 32 m and 128 m. As the condensate evolves in time, vortices are randomly creates at the borders of the tiles by the means of only wave function phase pattern (no time-dependent rotating potential). With a similar principle, in the second case we divide the wave function into pixels of the size 8 m (Pixel-8) or 4 m (Pixel-4) of the random phase each.
Appendix B links to animation
Here in the present section, we provide a list of links to movies of time evolution of the wave function , by means of the density field and phase field of (phase rescaled such that it is zero in the center ). The links accessed August 10th, 2025.
B.1 Conservative case and finite lifetime
Conservative case of long-living polaritons with ns:
Decaying polariton case with finite lifetime ns:
B.2 Simulations without high-energy wave damping ()
B.3 Simulations Tile-8 imprinting scheme for different tile grids
Appendix C Simulation snapshots for spoon stirring and Tile-8 at various time moments
Appendix D Simulations for Tile-8 imprinting Scheme and different tile block size
Appendix E Wave function at strong depletion